$A$ box contains $10$ good articles and $6$ with defects. One article is chosen at random. What is the probability that it is either good or has a defect?

  • A
    $0$
  • B
    $0.5$
  • C
    $0.75$
  • D
    $1$

Explore More

Similar Questions

For two events $A$ and $B$,if $P(A) = P(A|B) = 1/4$ and $P(B|A) = 1/2$,then

Find the probability that in a random arrangement of the letters of the word $DAUGHTER$,the letter $D$ occupies the first place.

In a simultaneous toss of $2$ coins,find the probability of getting exactly $1$ tail.

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

$A$ locker can be opened by dialing a fixed three-digit code (between $000$ and $999$). $A$ stranger who does not know the code tries to open the locker by dialing three digits at random. The probability that the stranger succeeds at the $k^{th}$ trial is

Difficult
View Solution

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo